LARENZIE:

(x+5)(2x−7) A x=4andx=7 B x=−5andx=72 C x=2andx=27 D x=5andx=72

LARENZIE:

help plz

seraphoftheend:

what kinda math is this like for what grade

LARENZIE:

i think its B

LARENZIE:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @seraphoftheend what kinda math is this like for what grade \(\color{#0cbb34}{\text{End of Quote}}\) wym

LARENZIE:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @LARENZIE \(\color{#0cbb34}{\text{Originally Posted by}}\) @seraphoftheend what kinda math is this like for what grade \(\color{#0cbb34}{\text{End of Quote}}\) wym \(\color{#0cbb34}{\text{End of Quote}}\) it's 11

LARENZIE:

yes I am i'm 17

LARENZIE:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shawnte Expand the polynomial using the FOIL method. \(\color{#0cbb34}{\text{End of Quote}}\) ok

snowflake0531:

uh no, we are trying to fidn the roots, not expand this

snowflake0531:

When you have two things multiplied together you can separate it (x+5)(2x−7)=0 We have x+5=0 and 2x-7 = 0 And then solve for the 2 x's

XLynx:

It's B. If you plug it into a graphing calculator, that's when it crosses the X axis.

LARENZIE:

thx boo